A rectangle and a square have the same area. The width of the rectangle is 6 cm less than the side of the square, and the length of the rectangle is 5 cm more than twice the side of the square. What are the dimensions of the rectangle?
Answer by ewatrrr(24785)

Area of rectangle(Lw) and Area of square (s^2) are the same:
s^2 = (2s+5)(s-6)
s^2 = 2s^2 - 7s - 30
s^2 - 7s - 30 = 0 (tossing out negative solution for unit measure)
(s+3)(s -10)= 0
s = 10cm
dimension of Rectangle: 25cm by 4cm
And...checking
10^2 = 25*4